• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Reitz

    Log entry time 00:03:55 on July22,2004

    Entry number 129197

    keyword=Shift Summary

    16:15 Arc energy measurement finished, Michelle starts testing
    her script to test BCM linearity
    17:00 tests done, moving to LH2 target
    17:10 production 3022 IHWP OUT
    18:00 MCC has problems with an IOC for the RF
    19:17 Beam back
    19:17 production 3023 IHWP OUT
    raster++ checkout looks weird, notice that BPMs are not
    updating -> when they switched to forced gain at 17:00
    the settings were wrong -> runs 3022/3023 are junk
    19:41 production 3024 IHWP OUT
    20:40 production 3025 IHWP OUT
    21:15 upper Coil of right septum reaches 7.8 K, we ask
    MCC to decrease beam to 36.5 (from 37.0)
    21:37 bleed through measurement -> 1% or 0.38uA
    21:43 production 3026 IHWP OUT
    22:30 Kent works on BCM linearity test,
    run 3027 IHWP OUT
    run 3028 IHWP OUT
    23:12 back to production, run 3029 IHWP OUT
    cont' into next shift